Q: Help to understand clearly and solve this example from Probability and Statistical Inference with the steps of the solution to better understand, thanks.

**Please give the step by step with details to completely see how the solution came about, plenty of thanks.

1) A random sample of 48 women who were test for cholesterol was classified according to age an cholesterol level and grouped into the following contingency table:

age <=210 >210 Totals
<50 16 8 24
>=50 7 17 24

Test the null hypothesis H_{0} that age and cholesterol level are independent attributes of classification of the 0.5% significance level.

What is the expected frequency value for cell 2,1?

a) 11.96

b) 14.04

c) 12.96

d) 11.50
Answer #1

Correct answer: option (D) 11.50

since

From the given data

<=210 >210 Total
<50 16 8 24
>=50 7 17 24
Total 23 25 48

E(2,1) = (Total of 2 row * total of 1 column)/ N = (24*23) /48 = 11.5